OpenStudy (anonymous):

the population of elephants, N, in a national park is modelled. a)the minimum number of elephants predicts please help!

OpenStudy (anonymous):

\[N=100+4t+400e ^{\frac{ -t }{ 10 }}\]

OpenStudy (cwrw238):

find the derivative : N' = 4 + 400 * (-1/10) * e(-t/10) = 0 at critical points solve this for t and substitute this value in the formula for N

OpenStudy (anonymous):

t=-loge(1/10) but the second question is asking em when this minimum occur when the park opened since 1 jan 1986 :( i am lost..

OpenStudy (cwrw238):

i get -10 ln (1/10) i'll re check

OpenStudy (anonymous):

yeh whats what i got as well but how could it be the minimum number of elephants predicted? and how do i know when this minimumm will occur?

OpenStudy (cwrw238):

-t/10 is the exponent right?

OpenStudy (cwrw238):

- 10 ln(1/10) is the time in years when the minimum is predicted what does the question as k for exactly?

OpenStudy (cwrw238):

that value ( time)is 23 .026 years

OpenStudy (anonymous):

what do you mean by k? there is no k on equation or on question..

OpenStudy (cwrw238):

lol - thats meant to be ask

OpenStudy (anonymous):

ah.. lol wait if it is asking me the minimum number of elephants predicted. shouldnt i sub in 0 into x?

OpenStudy (cwrw238):

no you should plug in t = -10 loge (1/10) into the formula for N = 100 + 4 (23.026) + 400*(e^(23.026)

OpenStudy (anonymous):

ohhh is that it? omg i am such a dumb

OpenStudy (cwrw238):

sorry i made a mistake in the last part its 100 + 4(23.026) + 400e^(-23.026/10)

OpenStudy (anonymous):

and then do i sub the number in to N again to know when it occur?

OpenStudy (cwrw238):

no when this occurs will be 1986 + 23.026 which will be in the year 2009

OpenStudy (cwrw238):

ok?

OpenStudy (cwrw238):

sorry - gotta go right now

OpenStudy (anonymous):

thank you anyways!

OpenStudy (anonymous):

btw what will be the date..

OpenStudy (anonymous):

@sirm3d can you help me?

OpenStudy (sirm3d):

did it say what t represents?

OpenStudy (anonymous):

nah but i calculated t=23.0259

OpenStudy (sirm3d):

t in years?

OpenStudy (anonymous):

yeh and it starts from jan 1st 1986. can you calculate the month and date as well?

OpenStudy (sirm3d):

assuming t = 23.059 years, convert that to years+days. count the number of days from Jan 1 1986 to get the approximate date (month, day, year).

OpenStudy (sirm3d):

0.059 yr x (365 days/1yr) = 21.535 days ~ 22 days that date would be Jan 23, 2009. i didn't check your solution if it was correct. i just assumed and showed you the computation for the nearest date the minimum would occur.

OpenStudy (anonymous):

thank you so much your help cleared my confusion! :)

OpenStudy (sirm3d):

yw. ^^